OpenStudy (anonymous):

If the three sides of a triangle are 4,5,6 then the cosine of the angle between the sides 4 and 15 is? Ans: A) 4/7 B) 5/7 C) 1/8 D) 3/16 E) 3/4 help me by showing the full process of the math..

OpenStudy (anonymous):

@abb0t @WhisperZ @Kreep @Superdeath @-BeYourself- any help here?

OpenStudy (anonymous):

@m&msdoodle can u help me here friend? :)

OpenStudy (anonymous):

i have no idea.... srry

OpenStudy (anonymous):

well... okay.... :(

OpenStudy (anonymous):

the triangle you described is  jf

OpenStudy (anonymous):

yes than??? @Garadon_Shagan

OpenStudy (anonymous):

Now you want angle a in terms of the sides 4 and 5 and cos. Consider side 4=C and side 5=B and 6=A The equation linking cos to an uneven triangle is A^2 = B^2 +C^2 - 2AC * cos a rearranging we have cos a = (A^2 - B^2 - C^2)/-2AC = (36 - 25 - 16)/(-2 * 6 * 4) = (-5)/(-48) = 5/48 which can't be simplified further which means I don't have a clue either. Sorry!

OpenStudy (anonymous):

I made a stupid mistake! note: -2BC cos a not BA! A^2 = B^2 + C^2 - 2BC cos a cos a = (-5)/(-2 * 5 * 4) cos a = 5/40 cos a = 1/8 meaning C is your answer

OpenStudy (anonymous):

@Garadon_Shagan thanks for the descriptive solution... i got it....